OpenStudy (anonymous):

If a city that currently has a population of 1000 triples in size every 8 years, what will the population be in 24 years and is the population growth modeled by a linear function or an exponential function? A) 10,000; linear B) 10,000; exponential C) 27,000; linear D) 27,000; exponential

OpenStudy (anonymous):

@baru please help

OpenStudy (baru):

initital population: 1000 can you calculate what the population is after 8 years ?

OpenStudy (baru):

@BrownieCat

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

3000

OpenStudy (anonymous):

@baru :)

OpenStudy (baru):

ok, whats it after 8 more years?

OpenStudy (anonymous):

d) because every 8 years it triples. At Beginning - Population 1000 After 8 years- Population = 3 * 1000 = 3000 After another 8 years (16 years from biginning) - Population = 3 * 3000 = 9000 After another 8 years (24 years from biginning) - Population = 3 * 9000 = 27000 It is plain to see that the growth is exponential since in is P = 1000*3^(years/8)